# Evacuation Planning on Time-Expanded Networks with Integrated Wildfire Information

### v1

@@ Abstract
We study the problem of evacuation planning for natural disasters, focusing on wildfire evacuations. By creating pre-planned evacuation routes that can be updated based on real-time data, we provide an easily adjustable approach to fire evacuation planning and implementation. Our method uses publicly available data and can be easily tailored for a particular region or circumstance.
We formulate large-scale evacuations as maximum flow problems on time-expanded networks, in which we integrate fire information given in the form of a shapefile. An initial flow is found based on a predicted fire, and is then updated based on revised fire information received during the evacuation.
We provide a proof on concept on three locations with historic deadly fires using data available through OpenStreetMaps, a basemap for a geographic information system (GIS), on a NetworkX Python script. The results validate viable running times and quality of information for application in practice.
@@
evacuation planning, maximum flow, time-expanded network
@@ 1 Introduction
Natural disasters pose a perpetual and significant threat to critical infrastructure and to human lives. Moreover, studies validate that the severity and frequency of natural disasters have intensified, with the greatest increases seen in recent decades [24, 30, 42]. Comprehensive preparation for these large-scale emergencies is, therefore, only becoming more imperative.
@@
Evacuation plans are a vital element of the planning process because research consistently demonstrates that successfully evacuating a disaster region is one of the most effective strategies in mitigating loss of life [23]. Furthermore, the lack of a well-defined evacuation plan can lead to appalling consequences; if an evacuation is not conducted efficiently, it often results in more loss [12, 14].
@@
One sobering illustration of the chaos that can occur took place in California in 2018. Butte County suffered the “Camp Fire”, which became the deadliest and most destructive wildfire in the state’s history [39, 54]. It resulted in 85 fatalities, in addition to destroying approximately 14,000 residences [31]. The fire started abruptly, causing major congestion on the roadways, and citizens abandoned their vehicles to continue the evacuation on foot. A key contributing factor was that the impacted area did not have sufficient road capacity on routes leaving the town. An investigation found that as the community had grown, infrastructure was unable to accommodate the population growth. This resulted in several roads with no exit, and only four functioning evacuation routes for the whole population [26].
@@
Wildfire evacuations like the one in Butte County, are an example of so-called “no-notice” evacuations. These occur when the disaster is unpredictable and requires rapid evacuation with little to no warning [8]. This is in contrast with “advanced-notice” evacuations, when there is enough time to warn the public about the disaster [18]. Evacuation plans are key in both instances, but are particularly important for no-notice evacuations [34]. In this case, authorities need to communicate precise instructions to the public as quickly and as clearly as possible, and therefore, must have an initial plan ready to go [60]. However, many evacuations, such as hurricanes, happen over a significant amount of time, meaning that it is also advantageous if plans can be modified based on the developing situation [15]. To construct an initial plan, models use past empirical data to predict how the disaster is most likely to unfold. However, during an emergency, if new data indicates something unexpected is occurring, the designated strategy must be updated as efficiently and quickly as possible. Throughout an evacuation, timing is critical, evacuate too early, and roads may become unnecessarily congested, hindering evacuations for other regions in the affected area. Evacuate too late and residents are left in danger [3]. Our method allows the user to judge the quality and feasibility of the initial evacuation plan in light of updated information and rapidly adjust when needed.
@@
The problem of finding optimal routes for no-notice evacuations has been well-studied, see [4, 41, 50] for detailed literature reviews. As in all mathematical modeling, evacuation models must balance the trade-off between complexity and detail. Small scale evacuations are able to take more detailed information into account such as human behavior, weather and landscape features, and traffic flow. However, when working with large road networks, it is not computationally feasible to model all these details. Model efficiency is absolutely critical if the hazard cannot be accurately predicted in advance. The ability to rapidly update an evacuation plan during an emergency is only possible if our model is solvable within a realistic time frame, in practice this means no more than a few minutes.
@@
In this work, we provide an easily adjustable algorithm that provides an initial community evacuation plan and then is able to update the plan based on current hazard conditions in real time. The information provided can be utilized by officials to quickly and properly direct evacuations, identify parts of the road network that will be used, and assist in the identification of bottlenecks. Our model takes into account factors such as decreased roadway capacity and the potential for blocked roads due to evolving conditions.
Our contribution focuses heavily on usability and adaptability. Because our package is freely available, self-contained, requires few computational resources, and uses open-source data, our algorithm can be used by experts and non-experts alike. The package itself outputs easy-to-read maps to be used in the field for real-time events. It also allows for various levels of detailed input from the user such as the tolerance and type of road network, alternative and/or additional hazard data, number of evacuees, and specific capacities of safe locations. Moreover, it can incorporate a range of different initial evacuation plans, assuming they are provided in the correct data type for the update. Even with this, we are able to model time down to the minute, create and update plans in approximately 1-2 minutes, and only need contract nodes in the road network that are less than 10 meters apart. That is to say, we achieve a viable balance between allowing for expert input, including fine details, and computational feasibility.
@@ 2.1 Dynamic Flows in Time-Expanded Networks
For computational purposes, it is often useful to transform a dynamic network into a time-expanded network: a single static network with copies of the node set for each time instance, and new arcs constructed from information based on the travel time of each arc in the dynamic network. In addition to being convenient, this removes the need to store arc travel time information explicitly. The nodes in the time-expanded network do not hold time-dependent information.
@@
$A^{M}:=\{(i(t),j(t^{\prime})):(i,j)\in A,\quad t^{\prime}=t+\lambda_{ij}(t)\leq T\}$
$A^{H}:=\{(i(t),i(t+1)):i\in(S\cup D),\quad t\in[T-1]_{0}\}.$
@@
$S_{T}=\{\ s_{i}(0)\in N_{T}:s_{i}\in S,s_{i}(0)=s_{i}\}$
$D_{T}=\{\ d_{j}(T)\in N_{T}:d_{j}\in D,d_{j}(T)=d_{j}+|N|\cdot T\}.$
Sources, $s_{i}(0)$, of the time-expanded network are all the copies of each source node, $s_{i}\in N$, from the original dynamic network that appear in the first time instance, $t=0$. The sink nodes of the time-expanded network are the final copy, $d_{j}(T)$, of each of the sink nodes, $d_{j}\in D$, of the dynamic network. These nodes are important to distinguish as they are the only nodes where we chose to allow holdover in our definition of a TEN. While the formulation could easily be adjusted for all nodes to have holdover, our definition encourages forward movement in the network. This is a desirable property given our specific application.
@@
We can now formally state the transformation as follows [16].
@@
$u_{i(t)j(t^{\prime})}$ $:=u_{ij}(t)$ for $\text{for }i,j\in A,t^{\prime}=t+\lambda_{ij}(t),t\in[T-1]_{0}$
$u_{i(t)i(t+1)}$ $:=a_{i}(t)$ for $\text{for }i\in N,t\in[T-1]_{0}.$
Using Definition 1, we can now reformulate the dynamic maximum flow problem on $G$, i.e., (LABEL:Dynamic), as a standard static multi-source, multi-sink maximum flow problem on the time-expanded network $G_{T}$ [29]:
@@ 2.2 Evacuation Models
Evacuation models can be categorized as either macroscopic or microscopic. Macroscopic models treat the crowd as a whole and represent their behavior as a flow rather than focusing on each person or vehicle as an individual. They use aggregate parameters such as density, flow rate, and average speed. In contrast, microscopic models consider individual behavior on a more detailed level, and can account for interactions between people or vehicles in the evacuation.
@@
The type of large-scale emergencies we consider, such as wildfires and tsunamis, necessitate the use of long-distance regional evacuation plans rather than small-scale plans, which are appropriate, for example, in a building fire [33]. Regional evacuations lead to the displacement of a large number of people using personal vehicles or other means of long-distance transportation, and due to the complexity, require the use of macroscopic evacuation models. In these large-scale events, several factors determine the outcome of an evacuation including warning time, response time, evacuation routes, and traffic flow [10, 35]
@@
Traffic assignment models [9] can be considered within the framework of User Equilibrium (UE) models, System Optimal (SO) models, Nearest Allocation (NA) models, and Constrained System Optimal (CSO) models. An evacuation plan is said to have User Equilibrium when no traveler can improve their individual travel time by changing routes (away from what the model prescribed) [53]. This desirable property may not be guaranteed when taking a purely System Optimal approach because here we maximize the total benefit of the system rather than considering the benefit to each individual separately. As such, some travelers may be assigned to routes that are much longer than the ones they would prefer to take. In contrast, when an evacuee uses a shortest path to reach their nearest shelter (based on geographical distance), this is considered a Nearest Allocation approach. Constrained System Optimal models seek the best of both worlds by combining SO with UE and NA to find an optimal solution for the system as a whole but given further constraints. The additional restrictions ensure that drivers are only assigned to “acceptable” or reasonable paths. For example, we could enforce that no driver should be assigned to a route that takes more than $x$ minutes longer than their shortest path to the given shelter. For a more detailed explanation of these properties in the context of evacuation planning see [49, 53, 56].
@@
Within these traffic modeling approaches, we can consider both the static and dynamic modeling of traffic flows [51]. In the case of static flow, we assume that the traffic flows are fixed for the entire planning horizon whereas a dynamic model allows flow rates to change over time. This is useful if we expect our hazard to impact the condition of our roads, for example. The common dynamic approaches are dynamic system optimal (DSO), dynamic user equilibrium (DUE), and dynamic traffic assignment (DTA).
@@
DTA models are optimization- and/or simulation-based models that are able to represent the time-varying characteristics of traffic conditions during an evacuation. Elements that have been considered within a DTA model framework for long-distance evacuations include maximum flow models [19], shortest path models [55], and cell-based transmission models [36]. Cell-based models (CTM) are a form of DTA which predict macroscopic traffic behavior by evaluating flow and density at a finite number of points and time steps along an arc. CTMs can be modeled as a linear program; however, due to the excessive number of constraints and variables for real-world networks, they require a prohibitive amount of memory and CPU time [27, 59]. Dynamic maximum flow models can also be modeled as linear programs and solved accordingly. However, they are more commonly solved using a specialized algorithm such as Ford-Fulkerson [16]. For an in-depth review of general DTA models, we refer the reader to [49].
@@
Most of the DTA models in the literature, such as [5, 8, 37, 43, 57], do not scale well and are not tractable for applications with large networks. A generally preferred methodology is to describe the static network at different points of time, i.e., to use a time-expanded version of the original network (TEN), see Definition 1 in Section 2.1. There are several papers which employ dynamic network flow models by constructing the TEN from the original static network [7, 21, 22, 32], and several versions of dynamic network flow problems are considered for evacuation planning. These include dynamic maximum flow, earliest arrival flow, quickest flow, and dynamic minimum cost flow [4].
@@
In [32], the authors model short-notice evacuations as a capacitated network flow problem on a dynamic time-expanded network. They develop an “evacuation scheduling algorithm” to obtain an evacuation schedule for each impacted region including prescribed evacuation paths and flow on the routes. The heuristic method regularly achieves high-quality solutions for even large networks on which exact optimal solutions are too computationally expensive to obtain. The model determines the number of people that can be evacuated and the order in which to send people to the safe locations they have been paired with. The granularity of the time interval used is 15 minutes. In this work, we build on several of the ideas from the evacuation literature presented here.

### v2

@@ Abstract
We study the problem of evacuation planning for natural disasters, focusing on wildfire evacuations. By creating pre-planned evacuation routes that can be updated based on real-time data, we provide an easily adjustable approach to evacuation planning and implementation. Our method uses publicly available data and can be tailored for a particular region or circumstance.
We formulate large-scale evacuations as maximum flow problems on time-expanded networks, in which we integrate hazard information given in the form of a shapefile. An initial flow and evacuation plan is found based on a predicted fire, and is then updated based on revised fire information received during the evacuation.
We provide a proof of concept on three locations with historic deadly fires using data available through OpenStreetMaps, a basemap for a geographic information system (GIS), on a NetworkX Python script. The results validate viable running times and quality of information for application in practice. Particular strengths are the scalability and modularity of our approach and accompanying software package.
@@
evacuation planning, maximum flow, time-expanded network, transportation, Geographical Information System, wildfire
@@ 1 Introduction
Natural disasters pose a persistent and significant threat to critical infrastructure and human lives. Moreover, studies validate that the severity and frequency of natural disasters have intensified, with the greatest increases seen in recent decades (Hooke, 2000; Kovács and Spens, 2009; Newkirk, 2001). Comprehensive preparation for these large-scale emergencies is, therefore, only becoming more imperative. Evacuation plans are a vital element of the planning process because research consistently demonstrates that successfully evacuating a disaster region is one of the most effective strategies in mitigating loss of life (Hobeika, 1984). Furthermore, the lack of a well-defined evacuation plan can lead to appalling consequences; if an evacuation is not conducted efficiently, it often results in more loss (Dixit and Wolshon, 2014; Federal Emergency Management Agency, 2019).
@@
One sobering illustration of the chaos that can occur took place in California in 2018. Butte County suffered the “Camp Fire”, which became the deadliest and most destructive wildfire in the state’s history (Maranghides et al., 2023; California Watershed Emergency Response Team, 2018). It resulted in 85 fatalities, in addition to destroying approximately 14,000 residences (Lam, 2019). The fire started abruptly, causing major congestion on the roadways, and citizens abandoned their vehicles to continue the evacuation on foot. A key contributing factor was that the impacted area did not have sufficient road capacity on routes leaving the town. An investigation found that as the community had grown, infrastructure was unable to accommodate the population growth. This resulted in several roads with no exit, and only four functioning evacuation routes for the whole population (John et al., 2018).
@@
Wildfire evacuations like the one in Butte County, are an example of so-called “no-notice” evacuations. These occur when the disaster is unpredictable and requires rapid evacuation with little or no warning (Chiu et al., 2007). This is in contrast with “advanced-notice” evacuations, when there is enough time to warn the public about the disaster (Golshani et al., 2019). Evacuation plans are key in both instances, but are particularly important for no-notice evacuations (Lindell et al., 2018). In this case, authorities need to communicate precise instructions to the public as quickly and as clearly as possible, and therefore, must have an initial plan ready to go (Zimmerman et al., 2007). However, many evacuations, such as hurricanes and wildfires, play out over a significant period of time, meaning that it is advantageous if plans can also be modified based on the developing situation (Federal Emergency Management Agency, 2021). To construct an initial plan, models can only use past empirical data to predict how the disaster is most likely to unfold. However, during an emergency, new data may indicate something unexpected, and then the designated strategy must be updated as efficiently and quickly as possible. Throughout an evacuation, timing is critical, evacuate too early, and roads may become unnecessarily congested, hindering evacuations for other regions in the affected area. Evacuate too late and residents are left in danger (Badiru et al., 2014). Our method allows the user to judge the quality and feasibility of the initial evacuation plan in light of updated information and rapidly adjust when needed.
@@
The problem of finding optimal routes for no-notice evacuations has been well-studied, see (Bayram, 2016; Murray-Tuite and Wolshon, 2013; Pel et al., 2012) for detailed literature reviews. As in all mathematical modeling, evacuation models must balance the trade-off between complexity and detail. Small scale evacuations are able to take more detailed information into account such as human behavior, weather and landscape features, and traffic flow. However, when working with large road networks, it is not computationally feasible to model all these details. Model efficiency is absolutely critical if the hazard cannot be accurately predicted in advance. The ability to rapidly update an evacuation plan during an emergency is only possible if our model is solvable within a realistic time frame, in practice this means no more than a few minutes.
@@
In this work, we provide an easily adjustable algorithm that provides an initial community evacuation plan and then is able to update the plan based on current hazard conditions in real time. The information provided can be utilized by officials to quickly and properly direct evacuations, identify parts of the road network that will be used, and assist in the identification of bottlenecks.
The fundamental component of our algorithm solves the maximum flow problem on a network (possibly with multiple source and sink nodes) that has been constructed to appropriately represent the geographical region of interest. Moreover, to account for the temporal aspect of an evacuation, we adapt this network to be so-called “time expanded” (Kotnyek, 2003), and further adjust the structure to account for changes caused by a developing hazard. Solutions to the maximum flow problem on a time-expanded network are already a critical tool in the evacuation planning literature, and our contribution builds on past work in a number of ways.
  • We provide a freely available and user-friendly self-contained software package, requiring minimal computational resources and using open-source data, which allows both experts and non-experts to produce and visualize evacuation plans for any geographic region covered by Open Street Maps (OpenStreetMap contributors, 2024).
  • The package may be used to produce initial evacuation plans for multiple and varying scenarios ahead of time which can be used to assist in emergency planning initiatives. Moreover, the model runs quickly enough that the user can update plans to account for any changes in real-time.
  • The package is highly-adaptable. For example, in our test instances, we model time to the minute and model the road network accurate to 10 meters ($\approx$ 30 feet). However, should the user require an alternative balance between fine detail and computational feasibility, these parameters can be easily changed.
  • Our package includes a procedure to model the development of a hazard over time, allowing us to take into account factors such as decreased roadway capacity and the potential for blocked roads due to evolving conditions. However, the modular format allows a user to instead easily incorporate data from more sophisticated models if this is available. Hence, our package can be used for accurate evacuation planning in multiple hazards and scenarios.
@@ 2.1 Dynamic Flows in Time-Expanded Networks
For computational purposes, it is often useful to transform a dynamic network into a time-expanded network: a single static network with copies of the node set for each time instance, and new arcs constructed from information based on the travel time of each arc in the dynamic network. In addition to being convenient, this removes the need to store arc travel time information explicitly. The nodes in the time-expanded network do not need to hold time-dependent information.
@@
$A^{M}:=\{(i(t),j(t^{\prime})):(i,j)\in A,\quad t^{\prime}=t+\lambda_{ij}(t)\leq T\}$ (7)
$A^{H}:=\{(i(t),i(t+1)):i\in(S\cup D),\quad t\in[T-1]_{0}\}.$ (8)
@@
$S_{T}=\{\ s_{i}(0)\in N_{T}:s_{i}\in S,s_{i}(0)=s_{i}\}$ (9)
$D_{T}=\{\ d_{j}(T)\in N_{T}:d_{j}\in D,d_{j}(T)=d_{j}+|N|\cdot T\}.$ (10)
Sources, $s_{i}(0)$, of the time-expanded network are all the copies of each source node, $s_{i}\in N$, from the original dynamic network that appear in the first time instance, $t=0$. The sink nodes of the time-expanded network are the final copy, $d_{j}(T)$, of each of the sink nodes, $d_{j}\in D$, of the dynamic network. These nodes are important to distinguish as they are the only nodes where we chose to allow holdover in our definition of a time-expanded network. The formulation could easily be adjusted for other nodes to have holdover.
@@
We can now formally state the transformation as follows.
@@
$u_{i(t)j(t^{\prime})}$ $:=u_{ij}(t)$ for $\text{for }i,j\in A,t^{\prime}=t+\lambda_{ij}(t),t\in[T-1]_{0}$ (12)
$u_{i(t)i(t+1)}$ $:=a_{i}(t)$ for $\text{for }i\in N,t\in[T-1]_{0}.$ (13)
Using Definition 1, we can now reformulate the dynamic maximum flow problem on $G$, i.e., (LABEL:Dynamic), as a standard static multi-source, multi-sink maximum flow problem on the time-expanded network $G_{T}$ (Kotnyek, 2003):
@@ 2.2 Evacuation Models
Evacuation models have been extensively studied, with the various approaches often categorized as either macroscopic or microscopic. Macroscopic models treat the crowd as a whole and represent their behavior as a flow rather than focusing on each person or vehicle as an individual. Macroscopic models address collective crowd behavior through traffic flow theories (Løvås, 1994), fluid dynamics (Henderson, 1974), and continuum mechanics (Hughes, 2002). They use aggregate parameters such as density, flow rate, and average speed. These approaches typically require modeling simplifications, such as converting the evacuation process to a target function (Bretschneider and Kimms, 2011) or representing the evacuation area as a network of links and nodes (Løvås, 1994).
In contrast, microscopic models consider individual behavior on a more detailed level, and can account for interactions between people or vehicles in the evacuation (Helbing and Molnár, 1995; Pan et al., 2006). These models employ various techniques including Social Force Models (Helbing and Molnár, 1995), Cellular Automata (Burstedde et al., 2001; Pelechano and Malkawi, 2008), Lattice Gas approaches (Helbing et al., 2003), and Agent-Based Models (Shi et al., 2009; Turner and Penn, 2002).
@@
The type of large-scale emergencies we consider, such as wildfires and tsunamis, necessitate the use of long-distance regional evacuation plans as opposed to small-scale plans, which are appropriate in a building fire, for example (Lindell, 2013). Regional evacuations lead to the displacement of a large number of people using personal vehicles or other means of long-distance transportation, and due to the complexity, require the use of macroscopic evacuation models.
Given this context, stage-based evacuation can be a useful approach. Here, emergency situations are managed by systematically prioritizing evacuation zones to minimize congestion and optimize resource utilization. The solution framework divides the affected network into zones based on their proximity to the hazard. This classification is important as zones closer to the incident location generally face shorter safety windows and experience more severe impacts (Liu et al., 2006a). While not directly addressed in this paper, our work can be easily extended for use within a stage-based model.
The literature on stage-based evacuation has contributed various methodological developments. For example, Liu et al. (2006a) presents a model that focuses on mitigating network congestion. This can be achieved in various ways depending on the specific instance, for example, having crowds exit through a limited number of gates (Lin et al., 2008), or reducing en-route wait time by adding delays to the departure announcement of different segments (Li et al., 2012). Other models prioritize the time aspect of evacuations by minimizing the total evacuation time (Chien and Korikanthimath, 2007), or by optimizing the warning time given to residents in varying evacuation stages (Chiu et al., 2007). Many specific contexts have also been studied, for example, Saadatseresht et al. (2009) and Stepanov and Smith (2009) consider multi-objective optimization within an urban setting. Objective functions considered include clearance time, total traveled distance, and blocking probability for evacuation planning. For additional models and information on stage-based evacuation, see (Li et al., 2012, 2018; Hosseini et al., 2021).
@@
In large-scale evacuations, such as those we consider in this work, several factors determine a successful outcome, including warning time, response time, evacuation routes, and how traffic flow is assigned to specific routes (Dash and Gladwin, 2007; Lindell and Prater, 2007). Traffic assignment models (Dafermos and Sparrow, 1969) can be considered within the framework of User Equilibrium (UE) models, System Optimal (SO) models, Nearest Allocation (NA) models, and Constrained System Optimal (CSO) models. An evacuation plan is said to have User Equilibrium when no traveler can improve their individual travel time by changing routes (away from what the model prescribed) (Sheffi, 1985). This desirable property may not be guaranteed when taking a purely System Optimal approach because here we maximize the total benefit of the system rather than considering the benefit to each individual separately. As such, some travelers may be assigned to routes that are much longer than the ones they would prefer to take. In contrast, when an evacuee uses a shortest path to reach their nearest shelter (based on geographical distance), this is considered a Nearest Allocation approach. Constrained System Optimal models seek the best of both worlds by combining SO with UE and NA to find an optimal solution for the system as a whole but given further constraints. The additional restrictions ensure that drivers are only assigned to “acceptable” or reasonable paths. For example, we could enforce that no driver should be assigned to a route that takes more than $x$ minutes longer than their shortest path to the given shelter. For a more detailed explanation of these properties in the context of evacuation planning see (Peeta and Ziliaskopoulos, 2001; Sheffi, 1985; Wardrop, 1952).
@@
Within these traffic modeling approaches, we can consider both the static and dynamic modeling of traffic flows (Transportation Research Board of the National Academies, 2010). In the case of static flow, we assume that the traffic flows are fixed for the entire planning horizon whereas a dynamic model allows flow rates to change over time. This is useful if we expect our hazard to impact the condition of our roads, for example. The common dynamic approaches are dynamic system optimal (DSO), dynamic user equilibrium (DUE), and dynamic traffic assignment (DTA).
@@
DTA models are optimization- and/or simulation-based models that are able to represent the time-varying characteristics of traffic conditions during an evacuation. Elements that have been considered within a DTA model framework for long-distance evacuations include maximum flow models (Guo et al., 2016), shortest path models (Wang et al., 2011), and cell-based transmission models (Liu et al., 2006c). Cell-based models (CTM) are a form of DTA which predict macroscopic traffic behavior by evaluating flow and density at a finite number of points and time steps along an arc. CTMs can be modeled as a linear program; however, due to the excessive number of constraints and variables for real-world networks, they require a prohibitive amount of memory and CPU time (Kimms and Maassen, 2011; Zheng and Chiu, 2011). Dynamic maximum flow models can also be modeled as linear programs and solved accordingly. However, they are more commonly solved using a specialized algorithm such as Ford-Fulkerson (Ford and Fulkerson, 1958). For an in-depth review of general DTA models, we refer the reader to (Peeta and Ziliaskopoulos, 2001).
@@
Most of the DTA models in the literature, such as (Bish et al., 2014; Chiu et al., 2007; Liu et al., 2006b; Ng and Waller, 2010; Yazici and Ozbay, 2010), do not scale well and are not tractable for applications with large networks. A generally preferred methodology is to describe the static network at different points of time, i.e., to use a time-expanded version of the original network (TEN), see Definition 1 in Section 2.1. There are several papers which employ dynamic network flow models by constructing the TEN from the original static network (Bretschneider and Kimms, 2011; Hamacher et al., 2013; Hamacher and Tjandra, 2002; Lim et al., 2012), and several versions of dynamic network flow problems are considered for evacuation planning. These include dynamic maximum flow, earliest arrival flow, quickest flow, and dynamic minimum cost flow (Bayram, 2016).
@@
In (Lim et al., 2012), the authors model short-notice evacuations as a capacitated network flow problem on a dynamic time-expanded network. They develop an “evacuation scheduling algorithm” to obtain an evacuation schedule for each impacted region including prescribed evacuation paths and flow on the routes. The heuristic method regularly achieves high-quality solutions for even large networks on which exact optimal solutions are too computationally expensive to obtain. The model determines the number of people that can be evacuated and the order in which to send people to the safe locations they have been paired with. The granularity of the time interval used is 15 minutes. In this work, we build on several of the ideas from the evacuation literature presented in this section.